Mr Haseeb has Rs 300,000,part at 9% and part at 8%.if total amount from two investment is Rs 25800,how much is invested at 9% and how much at 8%?
Answer by macston(5194)

N=amount at 9%; E=amount at 8%
N+E=300000
N=300000-E
0.08E+0.09N=25800 Substitute for N.
0.08E+0.09(300000-E)=25800
0.08E+27000-0.09E=25800 Subtract 27000 from each side.
-0.01E=-1200 Divide each side by 0.01.
E=120000 ANSWER 1: Rs 120000 was invested at 8%.
N=300000-E=300000-120000=180000 ANSWER 2: Rs 180000 was invested at 9%.
CHECK:
0.08E+0.09N=25800
0.08(120000)+0.09(180000)=25800
9600+16200=25800
25800=25800